OpenStudy (anonymous):

How many ways are there to choose a committee of 3 people from a group of 5 people? 10 12 60 70

OpenStudy (loser66):

\[\left(\begin{matrix}5\\3\end{matrix}\right)\]

OpenStudy (wolf1728):

The formula for this (choosing 'k' things from a larger set of 'n') is: n! / (k! * (n-k)!) 5! / (3! * 2!) = 5*4* 3! / 3! * 2 = 5 *4 / 2 =10

OpenStudy (anonymous):

Thank you!